### v2

# TimeBridge: Better Diffusion Prior Design with Bridge Models for Time Series Generation

Abstract.
Time series generation is widely used in real-world applications such as simulation, data augmentation, and hypothesis testing. Recently, diffusion models have emerged as the de facto approach to time series generation, enabling diverse synthesis scenarios. However, the fixed standard-Gaussian diffusion prior may be ill-suited for general time series data, such as temporal order and fixed points. In this paper, we propose TimeBridge, a framework that flexibly synthesizes time series data by using diffusion bridges to learn paths between a chosen prior and the data distribution. We then explore several prior designs tailored to time series synthesis. Our framework covers (i) data- and time-dependent priors for unconditional generation and (ii) scale-preserving priors for conditional generation. Experiments show that our framework with data-driven priors outperforms standard diffusion models on time series generation.
Time Series Generation, Diffusion Bridge, Diffusion Prior
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYccs: Computing methodologies Modeling methodologiesccs: Mathematics of computing Time series analysis
1. Introduction
Synthetic time series have expanded their applications by serving roles in data augmentation (Wen et al., 2021), data privacy (Wang et al., 2024), data compression (Zhou et al., 2023), and hypothesis‐driven simulations of data streams (Xia et al., 2024). Common tasks such as imputation and forecasting can be treated as special cases of time series generation. With advances in deep generative modeling, practitioners have developed VAE‐based (Naiman et al., 2024), GAN‐based (Yoon et al., 2019), and diffusion‐based methods (Ho et al., 2020; Karras et al., 2022). Diffusion approaches are especially effective, yielding high‐fidelity and diverse samples (Tashiro et al., 2021; Alcaraz and Strodthoff, 2023) by estimating the reverse trajectory of a forward noising process toward a standard Gaussian prior (Song et al., 2021).
Nevertheless, diffusion models may fail to capture characteristics of time series data because they rely on a fixed mapping from standard Gaussian noise to the data domain. Such restricted priors do not inherently reflect sequential order or scale (Kim et al., 2022), both of which are crucial for modeling historical statistics and temporal dependencies (Zeng et al., 2023). Adding domain‐specific constraints also remains difficult (Harvey et al., 2022; Du et al., 2023); for instance, Coletta et al. (2024) highlighted that generating data under specific bull or bear market patterns requires tailored embedding or penalty functions for sampling guidance.
Figure 1. Illustration of the proposed time series diffusion bridge framework enabling various prior selections.
Model Sampler Steps Uncond. Gen. Constrained Gen. Decomp. Layer Bridge Flexible Prior
[…]
Recent studies have revisited the noising process in diffusion models, exploring data-dependent forward and reverse distributions (Han et al., 2022; Yu et al., 2024) and joint consideration of diffusion steps with actual timestamps (Biloš et al., 2023; Chen et al., 2023a). We extend this direction through Schrödinger bridge (Schrödinger, 1932) and diffusion bridges (De Bortoli et al., 2021; Zhou et al., 2024), which learn optimal transport paths between the data distribution and a desired prior. By selecting appropriate fixed priors, diffusion bridges have demonstrated strong performance in conditional applications such as image-to-image translation (Zhou et al., 2024) and text-to-speech (Chen et al., 2023b).
In this paper, we propose TimeBridge, a flexible time series generation framework that addresses multiple time series synthesis with the prior design of diffusion bridge, as shown in Figure 1. Then, we ask what priors are best suited to time series generation as a replacement for the standard Gaussian prior, enabling more efficient and higher-quality synthesis across diverse settings. Against standard diffusion baselines compared in Table 1, the proposed TimeBridge outperforms in both unconditional and conditional tasks, achieving better results across several datasets in Figure 2.
Our key contributions with flexible prior selection are:
  • We explore diverse priors to alter the standard Gaussian for capturing time series characteristics and propose a new diffusion bridge model tailored to time series generation.
  • For unconditional generation, we design data- and time-dependent priors that reflect intrinsic temporal properties.
  • For conditional generation, we embed scale constraints into a pairwise prior and propose point-preserving sampling, extending its application to imputation tasks.
Figure 2. Performance measures (lower is better) on different tasks and sampling time (lower is better) of time series diffusion models. The proposed TimeBridge achieves both generalization in various tasks and sampling efficiency.
2. Preliminaries
Consider the multivariate time series ${\mathbf{x}}$ defined within the sample space $\mathcal{X}=\mathbb{R}^{d\times\tau}$, where $d$ represents the dimensionality of the variables and $\tau$ denotes the time length. For each index $i\in\{1,\ldots,N\}$ among the $N$ samples, each time series is denoted as ${\mathbf{x}}^{i}=({\bm{x}}^{i}(1),\ldots,{\bm{x}}^{i}(\tau))\in\mathcal{X}$, given the data ${\bm{x}}^{i}(k)\in\mathbb{R}^{d}$ for each timestamp $k\in\{1,\ldots,\tau\}$. For diffusion steps $t\in\{0,\ldots,T\}$, we denote intermediate time series samples as ${\mathbf{x}}_{t}$, and for a specific instance $i$ as ${\mathbf{x}}_{t}^{i}$.
2.1. Optimal Transport and Schrödinger Bridge
The objective of the Schrödinger Bridge (SB) problem is to find an entropy-regularized optimal transport path between the data distribution $p_{0}$ and a target prior $p_{T}$ by minimizing the KL divergence between paths (Léonard, 2012; Chen et al., 2021):
(1) $\underset{p\in\mathcal{P}_{[0,T]}}{min}D_{KL}(p\lVert p^{ref}),\quad s.t.\ p_{0}=p_{data},\ p_{T}=p_{prior},$
where, $\mathcal{P}_{[0,T]}$ denotes the space of path measures on the interval $[0,T]$, $p^{\text{ref}}$ is a reference path measure, and $p_{0},p_{T}$ are the marginals of $p$. By choosing a forward stochastic differential equation (SDE) associated with the reference process $p^{\text{ref}}$, the SB can be reformulated as following forward-backward SDEs:
(2) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g^{2}(t)\nabla\log\bm{\Psi}_{t}({\mathbf{x}}_{t})]dt+g(t)d{\mathbf{w}}_{t},\quad{\mathbf{x}}_{0}\sim p_{data},\end{split}$
(3) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g^{2}(t)\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},\quad{\mathbf{x}}_{T}\sim p_{prior},\end{split}$
where $\nabla\log\bm{\Psi}_{t}(x_{t})$ and $\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})$ are corresponding partial differential equations, detailed in Appendix A.
The Schrödinger bridge problem can be addressed using iterative methods (Kullback, 1968; Fortet, 1940) to simulate trajectories converging to the Schrödinger bridge problem (Vargas et al., 2021; Shi et al., 2024; De Bortoli et al., 2021). However, these iterative methods rely on expensive iterative approximation techniques and have seen limited empirical application (Zhou et al., 2024).
2.2. Diffusion Bridge
Standard Diffusion Models
In standard diffusion models (Ho et al., 2020), the diffusion process is constructed by gradually injecting noise into samples ${\mathbf{x}}_{0}$ drawn from the data distribution $p_{0}$, forwarding them into a fixed standard Gaussian distribution $p_{T}=\mathcal{N}(\bm{0},\bm{I})$. The corresponding forward SDE is:
(4) $d{\mathbf{x}}_{t}={\mathbf{f}}({\mathbf{x}}_{t},t)dt+g(t)d{\mathbf{w}}_{t},$
where ${\mathbf{f}}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ represents the drift function, $g:[0,T]\rightarrow\mathbb{R}$ denotes the diffusion coefficient, and ${\mathbf{w}}_{t}$ is the Wiener process. As diffusion models learn to reverse the Equation (4), the reverse SDE (Anderson, 1982) is formulated as follows:
(5) $d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},$
where $p({\mathbf{x}}_{t})$ refers to the probability density of ${\mathbf{x}}_{t}$. Diffusion models learn to match the score function $\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{t})$ for denoising (Song et al., 2021).
Diffusion Bridge
The aforementioned standard diffusion models cannot be extended to general prior distributions. Instead, we can construct the diffusion process towards any endpoint by applying Doob’s h-transform (Doob and Doob, 1984) to Equation (4). With a fixed endpoint ${\mathbf{y}}$, the transformed process for the diffusion bridge is as
(6) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g(t)^{2}{\mathbf{h}}({\mathbf{x}}_{t},t,{\mathbf{y}},T)]dt+g(t)d{\mathbf{w}}_{t},\\ {\mathbf{x}}_{0}\sim q_{data}({\mathbf{x}}),{\mathbf{x}}_{T}={\mathbf{y}},\end{split}$
where ${\mathbf{h}}({\mathbf{x}},t,{\mathbf{y}},T)=\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})|_{{\mathbf{x}}_{t}={\mathbf{x}},{\mathbf{x}}_{T}={\mathbf{y}}}$. For a given data distribution $q_{data}({\mathbf{x}},{\mathbf{y}})$, the diffusion bridge (Schrödinger, 1932; De Bortoli et al., 2021) aims to transport ${\mathbf{x}}_{0}$ to ${\mathbf{x}}_{T}$ where $({\mathbf{x}}_{0},{\mathbf{x}}_{T})=({\mathbf{x}},{\mathbf{y}})\sim q_{data}({\mathbf{x}},{\mathbf{y}})$ while following above forward process. Similar to the standard diffusion model, we can construct the reverse SDE as follows:
(7) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}({\mathbf{s}}({\mathbf{x}}_{t},t,{\mathbf{y}},T)-\\ {\mathbf{h}}({\mathbf{x}}_{t},t,{\mathbf{y}},T))]dt+g(t)d\bar{{\mathbf{w}}}_{t},\end{split}$
where ${\mathbf{x}}_{T}={\mathbf{y}}$ and its score function is calculated as
(8) ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)=\nabla_{{\mathbf{x}}_{t}}\text{log}q({\mathbf{x}}_{t}|{\mathbf{x}}_{T})|_{{\mathbf{x}}_{t}={\mathbf{x}},{\mathbf{x}}_{T}={\mathbf{y}}}.$
For training, diffusion bridge learns the score function ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$.
Diffusion bridge models have been investigated due to their ability to flexibly map in various domains, including image-to-image translation (Li et al., 2023; Zhou et al., 2024) and text-to-speech synthesis (Chen et al., 2023b), compared to standard diffusion models. Recently, adapting the advances of standard diffusion, such as consistency (He et al., 2024), and deterministic samplers (Zheng et al., 2025), has been further investigated in the diffusion bridge.
Note that although SB theory covers the concepts of diffusion bridges, we separate SB methods with iterative fitting and score-based diffusion bridge methods for clarity in this paper.
3. Framework
3.1. Scenarios
For time series generation, we cover both unconditional and conditional generations. Unconditional generation considers the total dataset containing $N$ samples, denoted as $\mathcal{D}=\{{{\mathbf{x}}^{i}}\}_{i=1}^{N}$. Our goal is to train a deep generative model to mimic the input dataset and yield synthetic samples statistically similar to a given time series dataset. We aim to make the distribution $p(\hat{\mathbf{x}}|\cdot)$ of synthetic data $\hat{\mathbf{x}}$, similar to the input data distribution $p({\mathbf{x}})$.
Conditional generation further utilizes the paired condition ${\mathbf{y}}$ to guide time series generation of each data sample ${\mathbf{x}}$ as $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. Thus, we try to make the conditional distribution $p(\hat{\mathbf{x}}|{\mathbf{y}})$ of synthetic data $\hat{\mathbf{x}}$ closely mirror the conditional distribution of input data $p({\mathbf{x}}|{\mathbf{y}})$. Considering that the condition in time series is often also a time series as ${\mathbf{y}}\in\mathcal{X}$, Coletta et al. (2024) divided the constraints into two types: (i) soft constraint to guide the time series generation such as trend, or (ii) hard constraint where we should follow during synthesis such as fixed points. In a broader view, the unmasked part in imputation or the historical data in forecasting can be considered as conditions for generation.
3.2. Diffusion Bridge for Time Series
To leverage the prior selections, we propose TimeBridge, a diffusion bridge tailored to time‐series synthesis. Our approach produces high-quality samples with fewer sampling steps and allows flexible prior selection. We now provide a detailed description of the sampler designs, network architecture, and overall framework.
Diffusion Bridge Sampler
To approximate the score ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$ in Equation (8), we adopt the Denoising Diffusion Bridge Model (DDBM) sampler (Zhou et al., 2024). Unlike prior time‐series methods based on the Schrödinger bridge (Chen et al., 2023b; Garg et al., 2024), DDBM aligns with recent diffusion frameworks and enables efficient bridge sampling. Accordingly, we exploit the tractable marginal distribution of ${\mathbf{x}}_{t}$ by setting ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\bm{\varepsilon}}$, where $\alpha_{t}$ and $\sigma_{t}$ are noise‐schedule functions and ${\bm{\varepsilon}}\sim\mathcal{N}(\bm{0},\bm{I})$. Under the variance‐preserving (VP) schedule, the marginal distribution at time $t$ becomes
(9) $q({\mathbf{x}}_{t}|{\mathbf{x}}_{0},{\mathbf{x}}_{T})=\mathcal{N}\left(\hat{\bm{\mu}}_{t},\hat{\sigma}^{2}_{t}\mathbf{I}\right),$
(10) $\hat{\bm{\mu}}_{t}=\textstyle\frac{SNR_{T}}{SNR_{t}}\frac{\alpha_{t}}{\alpha_{T}}{\mathbf{x}}_{T}+\alpha_{t}{\mathbf{x}}_{0}\left(1-\frac{SNR_{T}}{SNR_{t}}\right),$
(11) $\hat{\sigma}^{2}_{t}=\textstyle\sigma^{2}_{t}\left(1-\frac{SNR_{T}}{SNR_{t}}\right),$
where signal-to-ratio (SNR) is defined as $SNR_{t}={\alpha^{2}_{t}}/{\sigma^{2}_{t}}$. Since the mean $\hat{\bm{\mu}}_{t}$ is a linear combination of ${\mathbf{x}}_{0}$ and ${\mathbf{x}}_{T}$ at time $t$, the data scale is preserved for translation.
For score matching, Karras et al. (2022) proved that if a denoiser function $D({\mathbf{x}};\sigma)$ minimizes the expected error for samples drawn from $p_{data}$, then $D({\mathbf{x}};\sigma)$ becomes the linear function of score function, i.e., if $D({\mathbf{x}};\sigma)$ minimizes $\mathbb{E}_{{\mathbf{y}}\sim p_{data}}\mathbb{E}_{\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\bm{I})}\lVert D({\mathbf{y}}+\bm{{\bm{\varepsilon}}};\sigma)-{\mathbf{y}}\rVert^{2}_{2}$, then $\nabla_{\mathbf{x}}\text{log}p(x;\sigma)=(D({\mathbf{x}};\sigma)-{\mathbf{x}})/\sigma^{2}$. Thus, we match $D_{\theta}$ directly, which is equivalent to training the score function (Karras et al., 2022) as
(12) $\nabla_{{\mathbf{x}}_{t}}\text{log}q({\mathbf{x}}_{t}|{\mathbf{x}}_{T})\approx{\mathbf{s}}({D_{\theta}},{\mathbf{x}}_{t},t,{\mathbf{x}}_{T},T)$
$:=\frac{{\mathbf{x}}_{t}-(\frac{SNR_{T}}{SNR_{t}}\frac{\alpha_{t}}{\alpha_{T}}{\mathbf{x}}_{T}+\alpha_{t}D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})(1-\frac{SNR_{T}}{SNR_{t}}))}{\sigma^{2}_{t}(1-\frac{SNR_{T}}{SNR_{t}})},$
where $D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})$ is the model output as a denoiser.
Architecture Design
To obtain $D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})$, we follow the backbone of Diffusion-TS (Yuan and Qiao, 2024), which employs an encoder–decoder transformer. Let $w_{(\cdot)}^{i,t}$ denote the input of the interpretable layers for index $i\in{1,\ldots,K}$ at diffusion step $t$. Each decoder block is split into trend and seasonal synthesis layers. First, the trend component is synthesized with a polynomial regressor:
(13) $V_{tr}^{t}=\Sigma_{i=1}^{K}(C\cdot\text{Linear}(w_{tr}^{i,t})+X_{tr}^{i,t}),$
where $X_{tr}^{i,t}$ is the mean value of the output of the $i^{th}$ decoder block with the tensor multiplication ()($\cdot$)( ⋅ ) to $C$.
For seasonal components, we use the Fourier transformation as
(14) $S_{i,t}=\Sigma_{k=1}^{K}A_{i,t}^{\kappa_{i,t}^{(k)}}[cos(2\pi f_{\kappa_{i,t}^{(k)}}\tau c+\Phi_{i,t}^{\kappa_{i,t}^{(k)}})+cos(2\pi\bar{f}_{\kappa_{i,t}^{(k)}}\tau c+\bar{\Phi}_{i,t}^{\kappa_{i,t}^{(k)}})],$
where argTopK is to get the top $K$ amplitudes and $f_{k}$ represents the Fourier frequency of the index $k$ and $\bar{(\cdot)}$ denotes ()($\cdot$)( ⋅ ) of the conjugates. $A_{i,t}^{(k)}$ and $\Phi_{i,t}^{(k)}$ are the phase and amplitude of the $k$-th frequency, respectively, after the Fourier transform $\mathcal{F}$ as
(15) $A_{i,t}^{(k)}=\lvert\mathcal{F}(w_{seas}^{i,t})_{k}\rvert,\Phi_{i,t}^{(k)}=\phi(\mathcal{F}(w_{seas}^{i,t})_{k}),$
(16) $\kappa_{i,t}^{(1)},...,\kappa_{i,t}^{(K)}=\underset{k\in\{1,...,\lfloor\tau/2\rfloor+1\}}{\text{argTopK}}\{A_{i,t}^{(k)}\}.$
Overall Framework
Based on the previous subsections, we can design the output of the denoiser $D_{\theta}$ with both conditioning on noisy data ${\mathbf{x}}_{t}$ dependent on data ${\mathbf{x}}_{0}$ and prior ${\mathbf{x}}_{T}$ as follows:
(17) $D_{\theta}=V_{t}^{tr}(\theta,{\mathbf{x}}_{t},t,{\mathbf{x}}_{T})+\sum_{i=1}^{K}S_{i,t}(\theta,{\mathbf{x}}_{t},t,{\mathbf{x}}_{T})+R(\theta,{\mathbf{x}}_{t},t,{\mathbf{x}}_{T}),$
where $V_{t}^{tr}$ is the output for the trend synthesis layer, $S_{i,t}$ is the output for each seasonal synthesis layer $i$ among $K$ layers and $R$ is the output for estimated residual. Note that calculating Equation (17) is unique in diffusion bridge since the standard diffusion models (Yuan and Qiao, 2024) do not consider the dependency on prior ${\mathbf{x}}_{T}$ for calculating ${\mathbf{x}}_{t}$.
For training, our objective function $\mathcal{L}_{\theta}$ is formulated as follows:
(18) $\mathcal{L}_{\theta}=\mathbb{E}_{t,{\mathbf{x}}_{0}}[w_{t}(\lVert{\mathbf{x}}_{0}-D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})\rVert^{2}$
$+\lambda\lVert FFT({\mathbf{x}}_{0})-FFT(D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T}))\rVert^{2})].$
$w_{t}$ indicates the weight scheduler of the loss function, and $\lambda$ indicates the strength of the Fourier transform.
As time series data possess stochasticity (Shen and Kwok, 2023), we directly predict $D_{\theta}$ rather than using the reparametrization technique from noise prediction used in other domains (Karras et al., 2022; Zhou et al., 2024). Moreover, we utilize the second-order Heun sampler (Ascher and Petzold, 1998; Karras et al., 2022), achieving both quality and efficiency over first-order methods.
Overall, we name the proposed framework as TimeBridge, a flexible time series generation method that enables the selection of data-driven priors tailored to time series data based on (i) a flexible diffusion framework with diffusion bridge samplers and (ii) an interpretable decomposition-based time series structure.
4. Prior Design for Time Series Synthesis
Based on the flexibility of choosing priors with the proposed TimeBridge, we hypothesize that the standard Gaussian prior might not be the optimal choice. Thus, we investigate a general time series synthesis framework by analyzing the prior selection in diffusion models based on the following research questions (RQs):
RQ1. Are data-dependent priors effective for time series?
RQ2. How to model temporal dynamics with priors?
RQ3. Can constraints be used as priors for diffusion?
RQ4. How to preserve data points with diffusion bridge?
4.1. Data- and Time-dependent Priors for Time Series
For unconditional tasks, we examine the benefit of better priors depending on data and temporal properties of time series.
Data-Dependent Prior
To investigate RQ1, we suggest using data prior to having a data-dependent distribution to better approximate the data distribution. Sampling from $\mathcal{N}(\mathbf{0},\mathbf{I})$ lacks data-specific information, which can be enhanced by enforcing that the priors capture data scale and temporal dependencies. Recently, the usages of data-dependent prior have been investigated for audio (Popov et al., 2021; Lee et al., 2022) and image (Yu et al., 2024) domains. Therefore, we test the use of data-dependent prior for time series by setting the prior distribution as follows:
(19) ${\mathbf{x}}_{T}\sim\mathcal{N}(\bm{\mu},\texttt{diag}(\bm{\sigma}^{2})),$
where $\bm{\mu}$ and $\bm{\sigma}^{2}$ are the mean and variance of the data samples for each feature and timestamp independently calculated in the same dimension of the time series (e.g., for ETTh, resulting in $\bm{\mu}$ with a shape of $(24\times 7)$). $\texttt{diag}(\bm{\sigma}^{2})$ indicates the diagonal covariance matrix of each element.
Time-Dependent Prior
In response to RQ2, we focus on capturing temporal dependencies in noise addition towards prior distribution. Recently, Biloš et al. (2023) injected stochastic processes instead of random noise to preserve the continuity of temporal data and Han et al. (2022) utilized the pre-trained model for prediction and set the mean of prior as the output for regression.
As a solution, we employ Gaussian processes (GPs) for prior selection, which is effective for modeling temporal data (Biloš et al., 2023; Ansari et al., 2024). GP is a distribution over function that defines the joint variability of an arbitrary pair of time points. For given sequence of timestamps $\Omega=\{1,\ldots,\tau\}$ and a function $\bm{f}_{GP}$ of timestamps that follows a GP with mean function $\bm{m}(\cdot)$ and the covariance function $\bm{\Sigma}(\cdot)$, we can draw ${\mathbf{x}}_{T}$ as follows:
(20) $\bm{f}_{GP}\sim{GP}(\bm{m},\bm{\Sigma}),\ \ {\mathbf{x}}_{T}\sim\bm{f}_{GP}(\Omega).$
In this paper, we construct the time-dependent prior by aligning the GP with the aforementioned data-dependent prior at given time points. Specifically, we add the radial basis function (RBF) kernel $\mathcal{K}(i,j)=\eta exp(-\gamma|i-j|^{2})$ for timestamps $i$ and $j$ into Equation (19). Details for sampling method are in Appendix B.4.
4.2. Scale-preserving for Conditional Priors
We now consider the condition ${\mathbf{y}}\in\mathcal{X}$, provided as a time series, which is the most prevalent case for real-world settings. We explore using a prior on the same data scale instead of relying solely on conditional embeddings in standard diffusion.
Soft constraints with given trends
To answer RQ3, we argue that the diffusion bridge is well-suited for preserving a given condition by setting the prior as the condition. The diffusion bridge has shown its effectiveness in conditional tasks such as coloring (Zhou et al., 2024), image-to-image translation (Li et al., 2023; Liu et al., 2023), and improving naive base predictions (Chen et al., 2023b; Lee et al., 2024).
Conditional generation based on trend is straightforward by (i) setting the pair-wise condition to prior ${\mathbf{x}}_{T}={\mathbf{y}}$ and (ii) training the translation from trend to data. As the expected mean of ${\mathbf{x}}_{t}$ is a linear combination of ${\mathbf{x}}_{0}$ and ${\mathbf{x}}_{T}={\mathbf{y}}$ in Equation (9), the model learns the correction starting from the trend samples, utilizing the same data scale. In contrast, standard diffusion models are unsuitable for translating conditions into synthetic data; instead, they generate new data based on a specific trend condition, typically requiring additional steps such as providing guidance or correcting intermediate data samples ${\mathbf{x}}_{t}$ during sampling. Thus, we can also eliminate the need for penalty functions or corrections for sampling.
Hard constraints with fixed points
In RQ4, preserving data points of hard constraints can be important in conditional tasks. To ensure the fixed point condition during sampling, we introduce a novel point-preserving sampling method that prevents adding noise to the values to be fixed. This is a unique trait of our framework because standard diffusion models cannot preserve identity trajectories to the constrained values from Gaussian noise.
The direct case of fixed point constraints is imputation. With the mask ${\mathbf{m}}$ indicating the missing value as 0, we construct the condition ${\mathbf{c}}={\mathbf{y}}\odot{\mathbf{m}}$ for imputation where $\odot$ indicates the Hadamard product. However, unlike soft constraints that apply to all values in ${\mathbf{y}}\in\mathbb{R}^{\tau\times d}$, the condition for imputation ${\mathbf{c}}\in\mathbb{R}^{\tau\times d}$ has missing values, which should be interpolated to data-dependent values rather than zeros. Therefore, we build local linear models for condition-to-time series imputation. We adopt linear spline interpolation (De Boor, 1978) to generate a prior for fixed point constraints. In a feature-wise manner, let us consider $x\in\mathbb{R}$ as a single-dimensional input of ${\bm{x}}\in\mathbb{R}^{d}$. With a set $K$ of feature-wise observed times for $x$, the prior $x_{T}$ with the interpolation of each feature with timestamp $k$ is:
(21) $x_{T}(k)=x(k_{j-1})+\frac{x(k_{j})-x(k_{j-1})}{k_{j}-k_{j-1}}(k-k_{j-1}),$
$\quad k_{j-1}\leq k<k_{j},\quad j=1,\ldots,|K|-1.$
Note that before the first observation, $k<k_{0}$, $x_{T}(k)=x(k_{0})$. After the last observation, $k\geq k_{|K|-1}$, $x_{T}(k)=x(k_{|K|-1})$. We make details of the point-preserving sampler and its usage in imputation in Algorithm 1.
5. Related Works
Time Series Generation Models
Diffusion models show better performance in synthesis tasks, compared to time series generation models with VAE (Naiman et al., 2024), GANs (Yoon et al., 2019). Rasul et al. (2021) initially used diffusion networks for time series based on recurrent networks. For imputation, SSSD (Alcaraz and Strodthoff, 2023) and CSDI (Tashiro et al., 2021) considered time series imputation similar to image inpainting tasks. TimeDiff (Shen and Kwok, 2023), LDT (Feng et al., 2024), and TMDM (Li et al., 2024) focused on time series forecasting tasks. Coletta et al. (2024) investigated guiding the sampling with a penalty function called DiffTime. Our network design follows Diffusion-TS (Yuan and Qiao, 2024), which introduced a decomposition architecture to disentangle time series data into trend and seasonal components and applied a Fourier-based loss term. Recent studies also explore frequency-domain diffusion (Crabbé et al., 2024) and transformer-based alternatives (Chen et al., 2024).
Schrödinger Bridge in Time Series
Several studies have brought Schrödinger bridges to time series tasks. Chen et al. (2023a) analyzed algorithmic convergence and applied the bridge to imputation. Hamdouche et al. (2023) combined kernel-based bridges with feed-forward and LSTM networks for deep hedging. Garg et al. (2024) relaxed the Schrödinger bridge by setting a geometric mixture for the prior distribution and applied their method to time series. All of these methods require iterative solvers, which are computationally expensive and incompatible with standard diffusion pipelines. Also, None of them provides a unified framework for both unconditional and conditional generation. To the best of our knowledge, our paper is the first to use diffusion bridge, extending beyond classical Schrödinger bridge formulations and providing flexible priors.
Input: Model $D_{\theta}({\mathbf{x}}_{t})$, score function ${\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i}},t_{i},{\mathbf{x}}_{T},{T})$, diffusion steps $\{0=t_{0}<\cdots<t_{\Gamma}=T\}$, Condition ${\mathbf{y}}$, masking ${\mathbf{m}}$, step ratio $s$
Output: Denoised output ${\mathbf{x}}_{0}$
${\mathbf{x}}_{T}\leftarrow$ Set the prior as the hard constraint condition ${\mathbf{y}}$
if Imputation  then ${\mathbf{x}}_{T}\leftarrow$ Linear spline interpolation using Equation (21) with masking ${\mathbf{m}}$;
for $i=\Gamma,\ldots,1$ do
       Sample $\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0,I})$
       $\tilde{t}_{i}\leftarrow t_{i}+s(t_{i-1}-t_{i})$
       ${\mathbf{d}}_{t_{i}}\leftarrow-{\mathbf{f}}({\mathbf{x}}_{t_{i}},t_{i})+g^{2}(t_{i})({\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i}},t_{i},{\mathbf{x}}_{T},{T})-{\mathbf{h}}({\mathbf{x}}_{t_{i}},t_{i},{\mathbf{x}}_{T},T))$
       $\tilde{{\mathbf{x}}}_{t_{i}}\leftarrow{\mathbf{x}}_{t_{i}}+{\mathbf{d}}_{t_{i}}\odot{\mathbf{m}}(\tilde{t}_{i}-t_{i})+g(t_{i})\sqrt{\tilde{t}_{i}-t_{i}}\cdot\bm{{\bm{\varepsilon}}}\odot{\mathbf{m}}$
       $\tilde{{\mathbf{d}}}_{t_{i}}\leftarrow-{\mathbf{f}}(\tilde{{\mathbf{x}}}_{t_{i}},\tilde{t}_{i})+g^{2}(\tilde{t}_{i})(\frac{1}{2}{\mathbf{s}}(D_{\theta},\tilde{{\mathbf{x}}}_{t_{i}},\tilde{t}_{i},{\mathbf{x}}_{T},{T})-{\mathbf{h}}(\tilde{{\mathbf{x}}}_{t_{i}},\tilde{t}_{i},{\mathbf{x}}_{T},T))$
       ${\mathbf{x}}_{t_{i-1}}\leftarrow\tilde{{\mathbf{x}}}_{t_{i}}+\tilde{{\mathbf{d}}}_{t_{i}}\odot{\mathbf{m}}(t_{i-1}-\tilde{t}_{i})$
       if $i\neq 1$ then
             ${\mathbf{d}}_{t_{i}}^{\prime}\leftarrow-{\mathbf{f}}({\mathbf{x}}_{t_{i-1}},t_{i-1})+g^{2}(t_{i-1})(\frac{1}{2}{\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i-1}},t_{i-1},{\mathbf{x}}_{T},{T})-{\mathbf{h}}({\mathbf{x}}_{t_{i-1}},t_{i-1},{\mathbf{x}}_{T},T))$
            ${\mathbf{x}}_{t_{i-1}}\leftarrow\tilde{{\mathbf{x}}}_{t_{i}}+\frac{1}{2}({\mathbf{d}}_{t_{i}}^{\prime}+\tilde{{\mathbf{d}}}_{t_{i}})\odot{\mathbf{m}}(t_{i-1}-\tilde{t}_{i})$
Algorithm 1 Point-preserving Sampler for TimeBridge
Table 2. Quality comparison of unconditional time series synthetic data with length of 24. The best results are in bold and the second best are underlined. We report the average score (lower is better) and the rank of each method. The baseline experimental results (denoted by $\dagger$) are adopted from (Yuan and Qiao, 2024). We re-implement Diffusion-TS with their official code.
Metric Methods Synthetic datasets Real datasets Average (rank)
Sines MuJoCo ETTh Stocks Energy fMRI
Context-FID score (Lower is better) TimeVAE 0.307±.060 0.251±.015 0.805±.186 0.215±.035 1.631±.142 14.449±.969 2.943 (8)
TimeGAN 0.101±.014 0.563±.052 0.300±.013 0.103±.013 0.767±.103 1.292±.218 0.521 (6)
Cot-GAN 1.337±.068 1.094±.079 0.980±.071 0.408±.086 1.039±.028 7.813±.550 2.112 (7)
Diffwave 0.014±.002 0.393±.041 0.873±.061 0.232±.032 1.031±.131 0.244±.018 0.465 (5)
DiffTime 0.006±.001 0.188±.028 0.299±.044 0.236±.074 0.279±.045 0.340±.015 0.225 (4)
Diffusion-TS 0.008±.001 0.012±.002 0.129±.008 0.165±.043 0.095±.014 0.106±.007 0.086 (3)
TimeBridge 0.008±.001 0.018±.002 0.069±.004 0.079±.023 0.082±.007 0.097±.008 0.059 (2)
TimeBridge-GP 0.006±.001 0.016±.002 0.067±.007 0.054±.009 0.064±.007 0.096±.008 0.051 (1)
Correlational score (Lower is better) TimeVAE 0.131±.010 0.388±.041 0.111±020 0.095±.008 1.688±.226 17.296±.526 3.285 (6)
TimeGAN 0.045±.010 0.886±.039 0.210±.006 0.063±.005 4.010±.104 23.502±.039 4.786 (7)
Cot-GAN 0.049±.010 1.042±.007 0.249±.009 0.087±.004 3.164±.061 26.824±.449 5.236 (8)
Diffwave 0.022±.005 0.579±.018 0.175±.006 0.030±.020 5.001±.154 3.927±.049 1.622 (5)
DiffTime 0.017±.004 0.218±.031 0.067±.005 0.006±.002 1.158±.095 1.501±.048 0.495 (4)
Diffusion-TS 0.017±.006 0.193±.015 0.049±.012 0.011±.008 0.874±.163 1.151±.053 0.383 (3)
TimeBridge 0.024±.001 0.179±.011 0.035±.005 0.009±.009 1.064±.167 0.914±.028 0.371 (2)
TimeBridge-GP 0.018±.003 0.173±.025 0.034±.004 0.008±.008 0.902±.218 0.839±.018 0.329 (1)
Discriminative score (Lower is better) TimeVAE 0.041±.044 0.230±.102 0.209±.058 0.145±.120 0.499±.000 0.476±.044 0.267 (7)
TimeGAN 0.011±.008 0.238±.068 0.114±.055 0.102±.021 0.236±.012 0.484±.042 0.198 (5)
Cot-GAN 0.254±.137 0.426±.022 0.325±.099 0.230±.016 0.498±.002 0.492±.018 0.371 (8)
Diffwave 0.017±.008 0.203±.096 0.190±.008 0.232±.061 0.493±.004 0.402±.029 0.256 (6)
DiffTime 0.013±.006 0.154±.045 0.100±.007 0.097±.016 0.445±.004 0.245±.051 0.176 (4)
Diffusion-TS 0.008±.012 0.009±.011 0.074±.010 0.083±.044 0.126±.007 0.133±.024 0.072 (3)
TimeBridge 0.012±.003 0.007±.008 0.052±.004 0.052±.021 0.167±.003 0.077±.010 0.061 (2)
TimeBridge-GP 0.002±.002 0.006±.003 0.052±.002 0.049±.014 0.165±.009 0.082±.033 0.059 (1)
Table 3. Quality comparison of unconditional time-series synthetic data with a longer length of 64. Lower is better.
Dataset Method C.-FID ($\downarrow$) Corr. ($\downarrow$) Disc. ($\downarrow$)
ETTh Diffusion-TS 0.235±.011 0.061±.010 0.078±.017
TimeBridge-GP 0.110±.009 0.038±.011 0.045±.013
Stocks Diffusion-TS 0.948±.227 0.005±.005 0.149±.035
TimeBridge-GP 0.693±.165 0.005±.003 0.090±.018
Energy Diffusion-TS 0.067±.010 0.442±.077 0.065±.011
TimeBridge-GP 0.072±.010 0.548±.102 0.149±.005
fMRI Diffusion-TS 3.662±.104 4.585±.068 0.245±.234
TimeBridge-GP 0.527±.024 1.569±.028 0.211±.184
Average Diffusion-TS 1.228 1.273 0.134
TimeBridge-GP 0.351 0.540 0.124
6. Experiments
6.1. Experimental Setup
We assess the performance of TimeBridge using widely used time series datasets. We use two simulation datasets: Sines with 5 features of different frequencies and phases of sine functions and MuJoCo of multivariate advanced physics simulation with 14 features. For real-world datasets, we use ETT (Electricity Transformer Temperature) for long-term electric power with 7 features, Stocks of Google stock prices and volumes with 6 features, Energy, UCI dataset with appliances energy use in a low energy building with 28 features, and fMRI of the blood oxygen level-dependent functional MR imaging with 50 selected features.
For measures, we assess three metrics for unconditional time series generation on a normalized (0-1) scale. Context-Fréchet Inception Distance (Context-FID, C.-FID) score (Jeha et al., 2022) measures the distributional quality for mean and variance projected on embedding on trained TS2Vec (Yue et al., 2022). Correlational (Corr.) score (Liao et al., 2020) measures temporal dependency of real and synthetic data using the absolute error of cross-correlation. Discriminative (Disc.) score (Yoon et al., 2019) measures the performance of a classifier trained to distinguish original and synthetic data in a supervised manner. For imputation, we use the mean squared error (MSE) and mean absolute error (MAE).
Our experiments are conducted using PyTorch on four NVIDIA GeForce RTX 4090 GPUs. We report the mean and its interval of five experiments. For the detailed settings, refer to Appendix B.
Table 4. Quality comparison of trend-conditioned time series synthetic data.
Metric Methods ETTh Energy Average (rank)
Linear Polynomial Butterworth Linear Polynomial Butterworth
Context-FID score Trend Baseline 0.1729±.0228 0.1000±.0790 0.7343±.6795 0.8343±.0531 0.5864±.2016 0.8093±.2690 0.5395 (-)
SSSD 0.0137±.0028 0.0281±.0136 0.0575±.0104 0.0146±.0085 0.0135±.0015 0.0075±.0057 0.0225 (3)
Diffusion-TS 0.0049±.0008 0.0028±.0003 0.0030±.0003 0.0052±.0006 0.0050±.0005 0.0024±.0003 0.0039 (2)
TimeBridge 0.0028±.0003 0.0024±.0002 0.0023±.0002 0.0031±.0002 0.0036±.0002 0.0017±.0002 0.0027 (1)
Correlational score Trend Baseline 0.0493±.0104 0.0372±.0139 0.0524±.0194 0.4816±.0633 0.4816±.0999 0.5246±.0915 0.2711 (-)
SSSD 0.0362±.0137 0.0334±.0074 0.0404±.0032 0.6186±.1124 0.5594±.0925 0.5138±.0917 0.3003 (3)
Diffusion-TS 0.0247±.0109 0.0238±.0113 0.0239±.0127 0.4834±.2090 0.4843±.2116 0.4816±.2093 0.2536 (2)
TimeBridge 0.0240±.0120 0.0238±.0124 0.0237±.0123 0.4828±.2101 0.4830±.2106 0.4833±.2071 0.2534 (1)
Discriminative score Trend Baseline 0.1258±.0795 0.0574±.0626 0.1380±.0943 0.5000±.0633 0.4996±.0004 0.4997±.0003 0.3034 (-)
SSSD 0.0272±.0105 0.0301±.0142 0.1108±.0915 0.1871±.0749 0.2198±.0385 0.1360±.0797 0.1185 (3)
Diffusion-TS 0.0045±.0038 0.0038±.0040 0.0043±.0045 0.0031±.0035 0.0067±.0044 0.0046±.0034 0.0045 (2)
TimeBridge 0.0036±.0038 0.0023±.0021 0.0021±.0022 0.0054±.0026 0.0056±.0048 0.0037±.0035 0.0038 (1)
Table 5. Imputation results of MSE in the order of 1e-3 on the Mujoco dataset with a random mask of ratio {0.7, 0.8, 0.9}. The best results are in bold and the second best are underlined. The baseline experimental results are adopted from (Alcaraz and Strodthoff, 2023) and (Yuan and Qiao, 2024).
Methods
RNN
GRU-D
ODE
-RNN
Neural
CDE
Latent
-ODE
NAOMI NRTSI CSDI SSSD
Diffusion
-TS
TimeBridge
70% 11.34 9.86 8.35 3 1.46 0.63 0.24±.03 0.59±.08 0.37±.03 0.19±.00
80% 14.21 12.09 10.71 2.95 2.32 1.22 0.61±.10 1.00±.05 0.43±.03 0.26±.02
90% 19.68 16.47 13.52 3.6 4.42 4.06 4.84±.02 1.90±.03 0.73±.12 0.60±.02
Table 6. Imputation result of MSE and MAE on the ETTh and Energy datasets with a geometric mask of ratio {0.25, 0.5, 0.75}. The best results are in bold and the second best are underlined.
Metric MSE MAE
Methods SSSD Diffusion-TS TimeBridge TimeBridge-75% SSSD Diffusion-TS TimeBridge TimeBridge-75%
ETTh 25% 0.496±0.049 0.406±0.006 0.159±0.014 0.146±0.007 0.433±0.009 0.378±0.001 0.220±0.001 0.220±0.002
50% 0.523±0.061 0.561±0.010 0.226±0.010 0.199±0.006 0.435±0.013 0.422±0.001 0.248±0.002 0.237±0.002
75% 0.588±0.012 0.774±0.006 0.661±0.031 0.661±0.031 0.449±0.001 0.497±0.001 0.366±0.004 0.366±0.004
Energy 25% 108.82±19.12 137.77±1.14 17.06±2.34 12.03±0.40 1.94±0.19 1.39±0.00 0.47±0.01 0.45±0.00
50% 107.68±29.14 163.86±1.02 30.64±3.00 23.07±2.13 1.77±0.10 1.70±0.01 0.71±0.02 0.63±0.01
75% 139.07±4.49 199.38±1.82 63.84±2.15 63.84±2.15 1.87±0.01 2.09±0.01 1.07±0.01 1.07±0.01
6.2. Unconditional Generation
We demonstrate the unconditional generation tasks with a length of 24 as baselines. We compare with TimeVAE (Desai et al., 2021), TimeGAN (Yoon et al., 2019), Cot-GAN (Xu et al., 2020), Diffwave (Kong et al., 2021), DiffTime (Coletta et al., 2024) with SSSD (Alcaraz and Strodthoff, 2023) backbone, and Diffusion-TS (Yuan and Qiao, 2024). For the proposed TimeBridge, we choose the data-dependent settings in Equation (19) and TimeBridge-GP denotes the time-dependent GP prior in Equation (20). As the prior mean and variance are derived from the training datasets, the proposed method does not require additional information.
Results are shown in Table 2. We observe that the proposed method outperforms the previous methods. The performance gain is significant in Context-FID and discriminative scores, reducing the scores of 40.70% and 18.05% on average, respectively. Notably, the strength of data and time-dependent prior is shown in real datasets, such as ETTh and Stocks. This indicates that our prior selection and modeling help in approximating the data distribution during synthesis. In average rank, the proposed method shows the best performance. It indicates the power of prior selection in time series generation within a similar structure.
We compare our model with Diffusion-TS on long-horizon generation tasks with a sequence length of 64. The results in Table 3 show that, on average, TimeBridge-GP produces higher-quality samples than standard Gaussian. Additional experiments with even longer sequences are reported in Appendix C.
For visualization, we show the t-SNE plots with the generated data in Figure 3. The red original data and the blue synthetic data show an overlapped plot in almost all datasets. The generated samples well-approximate the distribution properties in complex datasets, such as the Stocks and Energy datasets.
(a) Sines
(b) Mujoco
(c) ETTh
(d) Stocks
(e) Energy
(f) fMRI
Figure 3. Visualization on t-SNE. Red and blue indicate original and synthetic data, respectively.
6.3. Synthesis with Trend Condition
We now consider trend-guided generation, where the data samples are in $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. For the trend ${\mathbf{y}}$, we consider three types: linear, polynomial with degree of three, and Butterworth filter (Butterworth et al., 1930), a signal processing filter that captures trends (Xia et al., 2024). Using these trends as conditions, we reimplement the experiments in Table 2 with the ETTh and Energy datasets. To apply the conditions to baseline methods, we feed the trend into a conditional embedding and concatenate it with the original embedding. For TimeBridge, we set the prior ${\mathbf{x}}_{T}={\mathbf{y}}$ upon the conditional embedding.
Experimental results are shown in Table 4. The rows with trend baseline indicate the results with the trends without any synthesis. After synthesis with trend conditions, we observe that the quality metrics diminished compared to the unconditional generation in Table 2, indicating that trend conditions help the model generate more realistic data samples. With the trend condition in ${\mathbf{x}}_{T}$, the proposed TimeBridge achieves the best performance in Context-FID in all settings and the best discriminative score in 5 out of 6 settings while maintaining correlation scores at similar levels to Diffusion-TS. This indicates that the proposed framework can be broadly adopted by different conditioning settings, without controlling any details other than the prior distribution. The sampling path of TimeBridge from trend to data is shown in Figure 4(a).
6.4. Imputation with Fixed Masks
For imputation, we compare the proposed method with imputation diffusion models and the conditional Langevin sampler of Diffusion-TS (Yuan and Qiao, 2024). We use random masks for the Mujoco dataset and the geometric mask from (Zerveas et al., 2021) for the ETTh and Energy datasets. We set the total length of imputation to 100 for the Mujoco dataset, matching the data length, and split the ETTh and Energy datasets into 48 time steps. For the proposed TimeBridge, we preserve the observed values and set the prior ${\mathbf{x}}_{T}$ as the interpolated values of the condition ${\mathbf{c}}$ as in Equation (21).
The imputation results are shown in Tables 5 and 6. Both tables demonstrate that TimeBridge achieves the lowest measures in imputation tasks. Since the Langevin sampler of Diffusion-TS does not require individual training for each missing ratio, we ensure a fair comparison by evaluating TimeBridge with the same training setting, TimeBridge-75%, which uses 75% missing values for training. TimeBridge-75% even outperforms the models trained with equal missing values, highlighting the generalization capability of TimeBridge in imputation tasks. The visualization of imputation results and sampling path can be found in Figures 5 and 4(b), respectively.
(a) Sampling of TimeBridge by applying the two pair-wise linear trends (red/blue) as priors.
(b) Imputation (red) of TimeBridge with point-preserving sampling using observed values (dots) of truth (green).
Figure 4. Illustration of sampling path from (Left) prior to (Right) data samples on the ETTh.
(a) ETTh Imputation
(b) Energy Imputation
Figure 5. Visualization of imputation on (a) ETTh and (b) Energy dataset. The green dotted lines and circles denote true and observed values with their intervals, respectively.
6.5. Additional Experiments
Ablation Study
Table 7 evaluates the effectiveness of training design with three variants: (1) variance exploding (VE) scheduler, (2) noise matching instead of directly predicting $D_{\theta}$ as in Zhou et al. (2024), and (3) without Fourier-based loss during training. The results indicate that each component of our diffusion scheduler and loss design contributes to improved performance.
Extended ablation results are provided in Appendix C.6.
Further Experiments
Please refer to Appendix C for further experimental results and findings, including the impact of prior selection measured by Wasserstein distance, evaluations on extreme data scenarios, and results of predictive score (Yoon et al., 2019).
6.6. Computation efficiency
Table 8 illustrates the strength of the proposed method, particularly when using the second-order Heun sampler during sampling. With fewer score function evaluations (NFE), our sampler achieves fast synthesis without compromising quality. We use 40 steps for sampling in all experiments, resulting in 119 NFEs with our settings. For comparison, we evaluate the Diffusion-TS using the same NFE. Given the trade-off of performance and sampling time in Diffusion-TS, our experimental results outperform previous methods. The marginal increase in time for TimeBridge is due to the time required to build the prior distribution.
Table 7. Ablation study for training and sampling features.
Metric Methods Datasets
ETTh Energy
Context-FID score (Lower is better) TimeBridge-GP 0.067±.007 0.064±.007
VE scheduler 0.256±.028 0.184±.003
F-matching 0.087±.010 0.081±.006
w/o Fourier 0.087±.010 0.093±.011
Discriminative score (Lower is better) TimeBridge-GP 0.052±.002 0.165±.009
VE scheduler 0.089±.009 0.236±.012
F-matching 0.052±.012 0.170±.007
w/o Fourier 0.060±.004 0.216±.007
Table 8. Comparison of sampling efficiency and generation performance. The time column indicates the generation time of each sample in milliseconds (ms) with a batch size of 512.
Data ETTh Stocks
Metric NFE Time C.-FID Disc. NFE Time C.-FID Disc.
Diffusion-TS 500 5.008 0.129 0.074 500 4.670 0.165 0.083
119 1.038 0.244 0.103 119 1.104 0.252 0.130
TimeBridge 119 1.426 0.067 0.052 119 1.429 0.079 0.052
7. Discussion and Conclusion
In this paper, we highlight the effectiveness of selecting prior distributions for time series generation using diffusion bridges. To address the limitations of the standard Gaussian prior in diffusion models, we propose a framework that allows the selection of any prior suitable for various tasks.
While the results are encouraging, several limitations remain. First, we rely on the standard architecture inherited from diffusion time-series synthesizers, which might not fully exploit the benefits of diffusion bridges. Future work should investigate transformer-based designs that incorporate prior information more effectively. Second, we leave the real-world applications with synthetic data beyond data synthesis itself for future study; realistic time-series synthesis could provide significant practical value.
As the proposed method can cover a wide range of scenarios in time series diffusion models, we believe this work significantly advances the investigation of prior distributions for developing a general time series generation framework. Ongoing advances in diffusion architectures will allow our approach to function as a plug-and-play adaptation that integrates with upcoming models.
Appendix A Diffusion Bridge
We introduce the details of the diffusion bridge mathematically, and how this framework is embraced in recent deep learning architectures.
A.1. Schrödinger Bridge
Standard diffusion models suffer from limitations such as the requirement of a sufficiently long time for the prior distribution to approximate a standard Gaussian distribution and the computational expense associated with sample generation. To address these challenges, diffusion Schrödinger bridge models have been recently introduced (De Bortoli et al., 2021; Chen et al., 2021). The goal of the Schrödinger bridge problem is to find an entropic optimal transport between two probability distributions in terms of Kullback–Leibler divergence on path spaces within a finite horizon (Chen et al., 2021; Léonard, 2012). While standard diffusion models and flow matching models are not guaranteed to provide optimal transport, Schrödinger bridge models aim to find paths that recover entropy-regularized versions of optimal transport (Shi et al., 2024). The formulation of the Schrödinger bridge problem is as follows:
(22) $\underset{p\in\mathcal{P}_{[0,T]}}{min}D_{KL}(p\lVert p^{ref}),\quad s.t.\ p_{0}=p_{data},\ p_{T}=p_{prior},$
where $\mathcal{P}_{[0,T]}$ refers to the space of path measures on $[0,T]$, $p^{ref}$ denotes the reference path measure, and $p_{0},p_{T}$ represents the marginal distributions of $p$ at each time step $0,T$, respectively. The diffusion bridge algorithm is introduced by utilizing $p^{ref}$ from the forward SDE in Equation (4) (Vargas et al., 2021; Chen et al., 2021; De Bortoli et al., 2021). Setting the reference path as the forward SDE, the problem becomes equivalent to following forward-backward SDEs:
(23) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g^{2}(t)\nabla\log\bm{\Psi}_{t}({\mathbf{x}}_{t})]dt+g(t)d{\mathbf{w}}_{t},\\ \quad{\mathbf{x}}_{0}\sim p_{data},\end{split}$
(24) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g^{2}(t)\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},\\ \quad{\mathbf{x}}_{T}\sim p_{prior},\end{split}$
where $\nabla\log\bm{\Psi}_{t}(x_{t})$ and $\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})$ are described by following PDEs:
(25) $\begin{split}\frac{\partial\bm{\Psi}}{\partial t}&=-\nabla_{x}\bm{\Psi}^{T}{\mathbf{f}}-\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\bm{\Psi})\\ \frac{\partial\hat{\bm{\Psi}}}{\partial t}&=-\nabla_{x}\cdot(\hat{\bm{\Psi}}{\mathbf{f}})+\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\hat{\bm{\Psi}})\\ &s.t.\quad\bm{\Psi}_{0}\hat{\bm{\Psi}}_{0}=p_{data},\bm{\Psi}_{T}\hat{\bm{\Psi}}_{T}=p_{prior},\\ &\qquad\quad p_{t}=\bm{\Psi}_{t}\hat{\bm{\Psi}}_{t}\ \text{for}\ t\in[0,T].\end{split}$
The Schrödinger bridge problem can be addressed using iterative methods (Kullback, 1968; Fortet, 1940). At the outset, several early studies employed iterative methods to simulate trajectories converging to the Schrödinger bridge problem (Vargas et al., 2021; Shi et al., 2024; De Bortoli et al., 2021). A classical solver for Schrödinger bridges is Iterative Proportional Fitting (IPF). Beginning with arbitrary half-potentials $(\bm{\Psi}^{(0)}_{t},\hat{\bm{\Psi}}^{(0)}_{t})$, the algorithm alternates two normalization steps:
[…]
After each full iteration, the product $\bm{\Psi}^{(n+1)}_{t}\hat{\bm{\Psi}}^{(n+1)}_{t}$ matches both marginals, the sequence converges geometrically, requiring one forward–backward integration per iteration. However, these methods rely on expensive iterative approximation techniques and have seen limited empirical application (Zhou et al., 2024).
A.2. Denoising Diffusion Bridge Models
Alternatively, Chen et al. (2021) leveraged the theory of forward-backward SDEs to derive an exact log-likelihood expression for the Schrödinger bridge, which accurately generalizes the approach for score generative models. Drawing inspiration from the perspective that the framework of diffusion bridges, employing forward-backward SDEs, encompasses the paradigms of score matching diffusion models (Song et al., 2020) and flow matching optimal transport paths (Lipman et al., 2022). Zhou et al. (2024) demonstrated that reintroducing several design choices from these domains becomes feasible. In particular, the reparameterization techniques outlined in Karras et al. (2022) are utilized.
Table 9. VP and VE instantiations of diffusion bridge settings in DDBM.
${\mathbf{f}}({\mathbf{x}}_{t},t)$ $g^{2}(t)$ $p({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ $SNR_{t}$ $\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})$
VP $\frac{d\text{log}\alpha_{t}}{dt}{\mathbf{x}}_{t}$ $\frac{d}{dt}\sigma^{2}_{t}-2\frac{d\text{log}\alpha_{t}}{dt}\sigma^{2}_{t}$ $\mathcal{N}(\alpha_{t}{\mathbf{x}}_{0},\sigma^{2}_{t}\mathbf{I})$ $\alpha^{2}_{t}/\sigma^{2}_{t}$ $\frac{(\alpha_{t}/\alpha_{T}){\mathbf{x}}_{T}-{\mathbf{x}}_{t}}{\sigma^{2}_{t}(SNR_{t}/SNR_{T}-1)}$
VE 0 $\frac{d}{dt}\sigma^{2}_{t}$ $\mathcal{N}({\mathbf{x}}_{0},\sigma^{2}_{t}\mathbf{I})$ $1/\sigma^{2}_{t}$ $\frac{{\mathbf{x}}_{T}-{\mathbf{x}}_{t}}{\sigma^{2}_{T}-\sigma^{2}_{t}}$
For noise scheduling, common options include variance-preserving (VP) and variance-exploding (VE) diffusion (Song et al., 2020). We express the reparameterization for the VP and VE bridge of Zhou et al. (2024) in Table 9. In our study, we choose to utilize the VP bridge, as our experimental findings indicate its superiority of VP over VE.
For score matching, Karras et al. (2022) proved that if a denoiser function $D({\mathbf{x}};\sigma)$ minimizes the expected $L_{2}$ error for samples drawn from $p_{data}$, then $D({\mathbf{x}};\sigma)$ becomes the linear function of score function, i.e., if $D({\mathbf{x}};\sigma)$ minimizes $\mathbb{E}_{{\mathbf{y}}\sim p_{data}}\mathbb{E}_{\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\bm{I})}\lVert D({\mathbf{y}}+\bm{{\bm{\varepsilon}}};\sigma)-{\mathbf{y}}\rVert^{2}_{2}$, then $\nabla_{\mathbf{x}}\text{log}p(x;\sigma)=(D({\mathbf{x}};\sigma)-{\mathbf{x}})/\sigma^{2}$. Thus, we can train $D_{\theta}$ directly instead of training the score function. In general, training a neural network to model $D_{\theta}$ directly is ineffective, due to its vast variance, which heavily depends on the noise level. Thus, Karras et al. (2022) reconstructed $D_{\theta}$ with $\sigma$-dependent skip connection as follows:
(28) $D_{\theta}({\mathbf{x}};\sigma)=c_{skip}(\sigma){\mathbf{x}}+c_{out}(\sigma)F_{\theta}(c_{in}(\sigma){\mathbf{x}};c_{noise}(\theta)),$
and trained $F_{\theta}$ instead of $D_{\theta}$ with the neural network. However, for time series data, estimating the diffusion noise poses greater challenges due to the presence of highly irregular noisy components (Shen and Kwok, 2023; Yuan and Qiao, 2024). Consistent with these studies, our experiments demonstrate that $D$-matching is more effective than $F$-matching. Detailed experimental results for comparison of VP with VE and $D$-matching with $F$-matching are provided in the ablation study in the Appendix C.6.
Appendix B Experimental Settings
B.1. Situations
We illustrate time series generation situations on various conditions, including unconditional, linear trend, polynomial trend, fixed point, and missing value, in Figure 6.
(a) Unconditional
(b) Linear trend
(c) Polynomial trend
(d) Fixed point
(e) Missing value
Figure 6. Illustration of time series generation situations under various conditions.
B.2. Baseline methods
For the unconditional baseline methods, we compare with TimeVAE (Desai et al., 2021), TimeGAN (Yoon et al., 2019), Cot-GAN (Xu et al., 2020), Diffwave (Kong et al., 2021), DiffTime (Coletta et al., 2024), and Diffusion-TS (Yuan and Qiao, 2024). For Table 2, we note the quality evaluation results in Diffusion-TS. We reimplement Diffusion-TS using their official GitHub in https://github.com/Y-debug-sys/Diffusion-TS, unfortunately, we cannot achieve the performance of Diffusion-TS in their original paper.
Furthermore, for the conditional generations, we reimplement SSSD (Alcaraz and Strodthoff, 2023) using the official GitHub https://github.com/SSSD/sssd. For the trend-guided synthesis, we treat the trend as conditional observations and train the SSSD model to match the total time series data. We add the conditional embedding layers in their original model architectures.
B.3. Evaluation Metrics
FID score is a widely used metric in image generation tasks, capturing the similarity between generated images and real ones (Heusel et al., 2017). To adapt it to the time series domain, Jeha et al. (2022) proposed Context-FID score, which replaces the Inception model used in the original FID with TS2Vec, a pre-trained time series representation learning model (Yue et al., 2022). As Context-FID score aligns with the predictive accuracy of the generative model in downstream tasks, a lower Context-FID score indicates higher synthetic data quality. For computation, first, we sample generated synthetic time series and real-time series separately. Next, both are encoded using a pre-trained TS2Vec model, and the FID score is calculated based on the resulting representations.
Correlational score is a metric used to evaluate the similarity in the covariance of time series features between original data and synthetic data (Liao et al., 2020). First, we estimate the covariance between the $i$-th and $j$-th features of the time series by:
(29) $cov_{i,j}=\frac{1}{T}\sum_{t=1}^{T}X_{i}^{t}X_{j}^{t}-\Big{(}\frac{1}{T}\sum_{t=1}^{T}X_{i}^{t}\Big{)}\Big{(}\frac{1}{T}\sum_{t=1}^{T}X_{j}^{t}\Big{)}.$
Then compute the correlation between the original data and synthetic data as follows:
(30) $\frac{1}{10}\sum_{i,j}^{d}\Big{|}\frac{cov_{i,j}^{o}}{\sqrt{cov_{i,i}^{o}cov_{j,j}^{o}}}-\frac{cov_{i,j}^{s}}{\sqrt{cov_{i,i}^{s}cov_{j,j}^{s}}}\Big{|},$
where $cov^{o}$ and $cov^{s}$ represent the covariance in the original data and synthetic data, respectively.
Discriminative score is a metric designed to evaluate the performance of a classifier in distinguishing between real and synthetic data. The score is calculated as the absolute difference between the classifier’s accuracy and 0.5.
B.4. Additional Experimental Details
Datasets
Table 10 presents the statistics of the datasets along with their online links.
Table 10. Details of datasets.
Dataset Sample size # of features Datatset link
Sines 10000 5 https://github.com/jsyoon0823/TimeGAN
Stocks 3773 6 https://finance.yahoo.com/quote/GOOG
ETTh 17420 7 https://github.com/zhouhaoyi/ETDataset
MuJoCo 10000 14 https://github.com/google-deepmind/dm_control
Energy 19711 28 https://archive.ics.uci.edu/dataset
fMRI 10000 50 https://www.fmrib.ox.ac.uk/datasets
Diffusion Sampler
We further explain the details of our sampler for the unconditional situation in Algorithm 2. For implementing the diffusion sampler, we modify the official code of DDBM (Zhou et al., 2024) in https://github.com/alexzhou907/DDBM as the backbone of our sampler.
Input: Diffusion model $D_{\theta}({\mathbf{x}}_{t})$, score function ${\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i}},t_{i},{\mathbf{x}}_{T},{T})$ defined in Equation 12, diffusion steps $\{0=t_{0}<\cdots<t_{\Gamma}=T\}$, step ratio $s$
Output: Denoised output ${\mathbf{x}}_{0}$
Sample ${\mathbf{x}}_{T}\sim\textit{prior}$ using Equations (19) or (20)
for $i=\Gamma,\ldots,1$ do
       Sample $\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
       $\tilde{t}_{i}\leftarrow t_{i}+s(t_{i-1}-t_{i})$
       ${\mathbf{d}}_{t_{i}}\leftarrow-{\mathbf{f}}({\mathbf{x}}_{t_{i}},t_{i})+g^{2}(t_{i})({\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i}},t_{i},{\mathbf{x}}_{T},{T})-{\mathbf{h}}({\mathbf{x}}_{t_{i}},t_{i},{\mathbf{x}}_{T},T))$
       $\tilde{{\mathbf{x}}}_{t_{i}}\leftarrow{\mathbf{x}}_{t_{i}}+{\mathbf{d}}_{t_{i}}(\tilde{t}_{i}-t_{i})+g(t_{i})\sqrt{\tilde{t}_{i}-t_{i}}\cdot\bm{{\bm{\varepsilon}}}$
       $\tilde{{\mathbf{d}}}_{t_{i}}\leftarrow-{\mathbf{f}}(\tilde{{\mathbf{x}}}_{\tilde{t}_{i}},\tilde{t}_{i})+g^{2}(\tilde{t}_{i})(\frac{1}{2}{\mathbf{s}}(D_{\theta},\tilde{{\mathbf{x}}}_{t_{i}},\tilde{t}_{i},{\mathbf{x}}_{T},{T})-{\mathbf{h}}(\tilde{{\mathbf{x}}}_{t_{i}},\tilde{t}_{i},{\mathbf{x}}_{T},T))$
       ${\mathbf{x}}_{t_{i-1}}\leftarrow\tilde{{\mathbf{x}}}_{t_{i}}+\tilde{{\mathbf{d}}}_{t_{i}}(t_{i-1}-\tilde{t}_{i})$
       if $i\neq 1$ then
             ${\mathbf{d}}^{\prime}_{t_{i}}\leftarrow-{\mathbf{f}}({\mathbf{x}}_{t_{i-1}},t_{i-1})+g^{2}(t_{i-1})(\frac{1}{2}{\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i-1}},t_{i-1},{\mathbf{x}}_{T},{T})-{\mathbf{h}}({\mathbf{x}}_{t_{i-1}},t_{i-1},{\mathbf{x}}_{T},T))$
            ${\mathbf{x}}_{t_{i-1}}\leftarrow\tilde{{\mathbf{x}}}_{t_{i}}+\frac{1}{2}({\mathbf{d}}^{\prime}_{t_{i}}+\tilde{{\mathbf{d}}}_{t_{i}})(t_{i-1}-\tilde{t}_{i})$
Algorithm 2 Unconditional Hybrid Sampler for TimeBridge
Hyperparameter Setups
For experimental setups, we follow the base setups of Yuan and Qiao (2024). Therefore, we try to use the minimal search space for determining only for diffusion scheduler ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\bm{\varepsilon}}$, where $\alpha_{t}$ and $\sigma_{t}$ are the functions of times.
As demonstrated in EDM (Karras et al., 2022), we can reparameterize $\alpha_{t}$ and $\sigma_{t}$ using $\beta_{min}$ and $\beta_{d}$ as follows:
(31) $\alpha_{t}$ $=e^{-\frac{1}{4}\beta_{d}t^{2}-\frac{1}{2}\beta_{min}t}$
(32) $\sigma_{t}$ $=e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}-1,$
With the aforementioned equations, we can adjust $\alpha_{t}$ and $\sigma_{t}$ based on $\beta_{min}$ and $\beta_{d}$.
Furthermore, we can adjust the weight scheduler of $w_{t}$ in Equation (18) with $\sigma_{data}$ as in EDM (Karras et al., 2022) and DDBM (Zhou et al., 2024) as follows:
(33) $w_{t}=\frac{(\sigma_{t}^{2}+\sigma_{data}^{2})}{(\sigma_{t}\cdot\sigma_{data})^{2}}.$
Note that EDM (Karras et al., 2022) and DDBM (Zhou et al., 2024) uses this $\sigma_{data}$ in the reparmetriziation of $F$-matching in Equation (28) to calculate, $c_{skip}$, $c_{out}$, $c_{in}$, $c_{noise}$, but we exclude them since we use direct $D$-matching.
In the GP setups for the unconditional generation, we additionally search the kernel parameter $\eta$.
The summarized hyperparameters used for unconditional generation and conditional generation are demonstrated in Tables 11 and 12 for unconditional and conditional settings, respectively.
Table 11. Hyperparameters for unconditional generation.
Setup Hyper-parameter Search space Sines MuJoCo ETTh Stocks Energy fMRI
Base setup (Yuan and Qiao, 2024) Batch size - 128 128 128 64 64 64
Training steps - 12000 14000 18000 10000 25000 15000
Attention heads / head dim - 4/16 4/16 4/16 4/16 4/24 4/24
Layers of encoder / decoder - 1/2 3/2 3/2 2/2 4/3 4/4
Warmup Learning rate - 0.008 0.008 0.008 0.008 0.008 0.008
Optimizer - Adam Adam Adam Adam Adam Adam
TimeBridge $\beta_{min}$0.1, 0.2} 0.2 0.1 0.2 0.2 0.2 0.1
$\beta_{d}$1, 2, 5, 10} 10 5 10 10 10 2
$\sigma_{data}$0.05, 0.1, 0.5} 0.5 0.5 0.1 0.1 0.05 0.5
TimeBridge-GP Kernel parameter $\eta$0.1, 0.3, 0.5, 1} 1 0.1 0.5 0.3 1 0.1
Table 12. Hyperparameters for conditional generation. $\dagger$ We use 0.025 for the Energy dataset with Butterworth trend setting.
Method Hyper-parameter Search space Trend & Imputation
Mujoco ETTh Energy
TimeBridge $\beta_{min}$0.1} 0.1 0.1 0.1
$\beta_{d}$0.2, 0.5, 1} 0.2 0.2 0.2
$\sigma_{data}$0.05, 0.1, 0.5} 0.5 0.1 0.05
We unified the sampling settings with the second-order Heun sampler with 40 steps, which have 119 NFE. We use the churn ratio $s=0.33$ for stochastic sampling in Algorithms 1 and 2.
Details for sampling from GP Prior
We sample ${\mathbf{x}}_{T}$ in Equation (20) as follows:
(34) ${\mathbf{x}}_{T}\sim\mathcal{N}(\bm{m}(\Omega),\bm{\Sigma}(\Omega)),\quad\Omega=\{1,\ldots,\tau\}.$
Specifically, we set $\bm{m}(\Omega)$ and $\bm{\Sigma}(\Omega)$ by $\bm{m}(\Omega)=\bm{\mu}$ and $\bm{\Sigma}(\Omega)_{i,j}=\texttt{diag}(\bm{\sigma}^{2})_{i,j}+\exp\left(-\frac{|i-j|^{2}}{2l^{2}}\right)$ for timestamps $i$ and $j$ in Equation (19). $l$ is the normalization parameter using data standard deviation to encode data information in kernel function.
Appendix C Additional Experimental Results
C.1. Wasserstein distance
To experimentally justify the strength of the diffusion prior, we measure the Wasserstein distance within the diffusion bridge’s optimal transport framework. Kollovieh et al. (2024) argue that closer alignment of data and prior distributions reduces path complexity, and time-reflective priors can lower the Wasserstein distance. As shown in Table 13, our data-dependent priors (unconditional and imputation) yield shorter distances. While not always a direct indicator of performance, this may explain why diffusion bridge models surpass standard diffusion approaches.
Table 13. Wasserstein distance results for unconditional and imputation settings using different priors.
Unconditional Imputation
Prior $\mathcal{N}(\mathbf{0},\mathbf{I})$ $\mathcal{N}(\bm{\mu},\bm{\Sigma})$ (Ours) $\mathcal{N}(\mathbf{0},\mathbf{I})$ Linear Spline (Ours)
ETTh 14.285 2.387 8.000 0.485
Energy 28.128 6.376 16.113 1.494
C.2. Extreme cases
To evaluate effectiveness on extreme values, we perform an ablation study on tailed data samples using real-world data in our experiments, i.e., ETTh, Stocks, and Energy datasets. For each real and synthetic dataset, we extract top $k\%$ data samples with high variance as extreme data, respectively. The results from samples with the top $20\%$ and $50\%$ variance are shown in Table 14. Our method outperformed 10 out of 12 metrics, showing its robustness in high variance and tailed data.
Table 14. Ablation results on extreme values of top $k\%$ variance.
Top $k\%$ Metric ETTh Energy Stock
Diffusion-TS Ours Diffusion-TS Ours Diffusion-TS Ours
20% C.-FID 0.153 0.103 0.128 0.093 0.788 0.257
Corr. 0.095 0.070 0.946 0.150 1.217 0.010
50% C.-FID 0.144 0.098 0.139 0.096 0.475 0.154
Corr. 0.062 0.048 0.886 0.021 1.028 0.012
C.3. Longer Sequences
Modeling longer sequences is crucial for time series generation. Thus, we have tested lengths of 192 and 360, common for forecasting cyclic or annual patterns. By controlling only $sigma_{d}ata$ and fixing other parameters to minimize the search space, Table 15 shows that our method outperforms standard diffusion priors, excelling in 11 out of 12 metrics. These results will be included in the main paper for real-world usage.
Table 15. Comparison of longer sequence lengths (192, 360). Lower is better.
Context-FID Cross-Correlation Discriminative
Dataset (length) Diffusion-TS Ours Diffusion-TS Ours Diffusion-TS Ours
Stocks (192) 0.679 0.548 0.009 0.006 0.144 0.060
Stocks (360) 0.954 0.606 0.016 0.004 0.125 0.069
ETTh (192) 1.021 1.535 0.106 0.098 0.187 0.096
ETTh (360) 1.977 1.780 0.114 0.099 0.226 0.109
C.4. Generation with same NFE
Table 16 presents the experimental results for Diffusion-TS (Yuan and Qiao, 2024) with the same NFE of 119 across most datasets. The results demonstrate that TimeBridge outperforms Diffusion-TS by using better priors for time series and the second-order sampler, compared to standard Gaussian priors.
Table 16. Comparison of generation performance with same NFE of 119.
Metric Methods Synthetic datasets Real datasets
Sines MuJoCo ETTh Stocks Energy fMRI
Context-FID score (Lower is better) Diffusion-TS (Yuan and Qiao, 2024) 0.014±.001 0.033±.006 0.244±.012 0.252±.014 0.126±.026 0.149±.009
TimeBridge-GP 0.006±.001 0.016±.002 0.067±.007 0.054±.009 0.064±.007 0.096±.008
Correlational score (Lower is better) Diffusion-TS (Yuan and Qiao, 2024) 0.015±.002 0.194±.023 0.071±.011 0.021±.004 0.878±.144 1.079±.024
TimeBridge-GP 0.018±.003 0.173±.025 0.034±.004 0.008±.008 0.902±.218 0.839±.018
Discriminative score (Lower is better) Diffusion-TS (Yuan and Qiao, 2024) 0.043±.014 0.040±.008 0.103±.006 0.130±.009 0.143±.008 0.156±.016
TimeBridge-GP 0.002±.002 0.006±.003 0.052±.002 0.049±.014 0.165±.009 0.082±.033
Predictive score (Lower is better) Diffusion-TS (Yuan and Qiao, 2024) 0.095±.000 0.008±.001 0.125±.005 0.038±.000 0.251±.000 0.099±.000
TimeBridge-GP 0.093±.000 0.009±.001 0.121±.006 0.037±.000 0.251±.000 0.100±.000
C.5. Predictive score
Predictive score (Yoon et al., 2019) measures the performance to predict the next step trained with synthetic data as a train-synthesis-and-test-real (TSTR) method. Table 17 illustrates the predictive score in the same experimental setting of Table 2. Similarly, Table 18 shows the predictive score of the same experiments as in Table 4.
Table 17. Predictive score unconditional time series synthetic data. The best results are in bold and the second best are underlined. The baseline experimental results (denoted by $\dagger$) are adopted from (Yuan and Qiao, 2024). We re-implement Diffusion-TS with their official code.
Metric Methods Synthetic datasets Real datasets
Sines MuJoCo ETTh Stocks Energy fMRI
Predictive score (Lower is better) TimeVAE (Desai et al., 2021) 0.093±.000 0.012±.002 0.126±.004 0.039±.000 0.292±.000 0.113±.003
TimeGAN (Yoon et al., 2019) 0.093±.019 0.025±.003 0.124±.001 0.038±.001 0.273±.004 0.126±.002
Cot-GAN (Xu et al., 2020) 0.100±.000 0.068±.009 0.129±.000 0.047±.001 0.259±.000 0.185±.003
Diffwave (Kong et al., 2021) 0.093±.000 0.013±.000 0.130±.001 0.047±.000 0.251±.000 0.101±.000
Diffusion-TS (Yuan and Qiao, 2024) 0.093±.000 0.007±.001 0.123±.003 0.037±.000 0.251±.000 0.100±.000
TimeBridge 0.093±.000 0.009±.001 0.121±.006 0.037±.000 0.251±.000 0.100±.000
TimeBridge-GP 0.093±.000 0.009±.001 0.121±.006 0.037±.000 0.251±.000 0.100±.000
Table 18. Quality comparison of trend-conditioned time series synthetic data.
Metric Methods ETTh Energy
Linear Polynomial Butterworth Linear Polynomial Butterworth
Predictive score Trend Baseline 0.1208±.0035 0.1200±.0016 0.1210±.0030 0.3684±.0085 0.4065±.0362 0.3744±.0413
SSSD (Alcaraz and Strodthoff, 2023) 0.1228±.0041 0.1206±.0036 0.1214±.0047 0.2498±.0005 0.2501±.0004 0.2498±.0004
Diffusion-TS (Yuan and Qiao, 2024) 0.1204±.0055 0.1200±.0059 0.1197±.0060 0.2503±.0003 0.2500±.0005 0.2495±.0007
TimeBridge 0.1195±.0062 0.1196±.0061 0.1196±.0060 0.2501±.0002 0.2500±.0005 0.2502±.0003
C.6. Ablation study
VP and VE scheduler
We perform experiments to compare the performance of the VP bridge and VE bridge for unconditional generation. The results, as shown in Table 19, indicate that the VP bridge consistently outperforms the VE bridge in all measures. We hypothesis that the enlarging level of noise addition in VE might be harmful to time series generation.
Table 19. Quality comparison of unconditional time series with VP and VE scheduler.
Metric Methods Synthetic datasets Real datasets
Sines MuJoCo ETTh Stocks Energy fMRI
Context-FID score (Lower is better) TimeBridge-GP-VP 0.006±.001 0.016±.002 0.067±.007 0.054±.009 0.064±.007 0.096±.008
TimeBridge-GP-VE 0.011±.002 0.031±.004 0.256±.028 0.130±.028 0.184±0.03 0.102±.009
Correlational score (Lower is better) TimeBridge-GP-VP 0.018±.003 0.173±.025 0.034±.004 0.008±.008 0.902±.218 0.839±.018
TimeBridge-GP-VE 0.027±.003 0.190±.010 0.060±.011 0.065±.006 1.115±.160 0.895±.021
Discriminative score (Lower is better) TimeBridge-GP-VP 0.002±.002 0.006±.003 0.052±.002 0.049±.014 0.165±.009 0.082±.033
TimeBridge-GP-VE 0.017±.009 0.026±.011 0.089±.009 0.095±.000 0.236±.012 0.109±.013
Predictive score (Lower is better) TimeBridge-GP-VP 0.093±.000 0.009±.001 0.121±.006 0.037±.000 0.251±.000 0.100±.000
TimeBridge-GP-VE 0.093±.000 0.010±.001 0.122±.005 0.038±.000 0.251±.000 0.100±.000
Importance of $D$-matching in time series
To verify the effectiveness of $D$-matching in time series, we compare the performance of $D$-matching and $F$-matching methods for unconditional generation. The experimental results are shown in Table 20. It can be observed that $D$-matching yields better results in most metrics across the majority of datasets in our experiments.
Table 20. Quality comparison of unconditional time series with $D$-matching and $F$-matching.
Metric Methods Synthetic datasets Real datasets
Sines MuJoCo ETTh Stocks Energy fMRI
Context-FID score (Lower is better) TimeBridge-D 0.008±.001 0.018±.002 0.069±.004 0.079±.023 0.082±.007 0.097±.008
TimeBridge-F 0.011±.002 0.023±.003 0.093±.010 0.060±.013 0.101±.006 0.104±.008
TimeBridge-GP-D 0.006±.001 0.016±.002 0.067±.007 0.054±.009 0.064±.007 0.096±.008
TimeBridge-GP-F 0.019±.004 0.019±.004 0.087±.010 0.068±.014 0.081±.006 0.106±.008
Correlational score (Lower is better) TimeBridge-D 0.024±.001 0.179±.011 0.035±.005 0.009±.009 1.064±.167 0.914±.028
TimeBridge-F 0.016±.002 0.184±.017 0.037±.004 0.011±.007 0.904±.134 1.076±.048
TimeBridge-GP-D 0.018±.003 0.173±.025 0.034±.004 0.008±.008 0.902±.218 0.839±.018
TimeBridge-GP-F 0.033±.009 0.181±.028 0.038±.006 0.007±.005 0.885±.246 1.104±.031
Discriminative score (Lower is better) TimeBridge-D 0.012±.003 0.007±.008 0.052±.004 0.052±.021 0.167±.003 0.077±.010
TimeBridge-F 0.031±.013 0.016±.011 0.053±.017 0.070±.011 0.167±.006 0.045±.026
TimeBridge-GP-D 0.002±.002 0.006±.003 0.052±.002 0.049±.014 0.165±.009 0.082±.033
TimeBridge-GP-F 0.006±.005 0.019±.010 0.052±.012 0.065±.017 0.170±.007 0.036±.038
Predictive score (Lower is better) TimeBridge-D 0.093±.000 0.009±.001 0.121±.006 0.037±.000 0.251±.000 0.100±.000
TimeBridge-F 0.095±.000 0.008±.001 0.121±.006 0.037±.000 0.251±.000 0.100±.000
TimeBridge-GP-D 0.093±.000 0.009±.001 0.121±.006 0.037±.000 0.251±.000 0.100±.000
TimeBridge-GP-F 0.093±.000 0.008±.001 0.121±.006 0.037±.000 0.251±.000 0.100±.000
Importance of stochasticity
We conduct experiments to compare the performance of the SDE solver with and without the stochasticity during sampling. By setting the step ratio to $s=0$, we prevent noise from being added during sampling. The experimental results are presented in Table 21. The results show that the SDE solver performs better than the non-stochastic method. Note that TimeBridge-GP uses $s=0.33$.
Table 21. Quality comparison of unconditional time series with sampling step ratio $s$.
Metric Methods Synthetic datasets Real datasets
Sines MuJoCo ETTh Stocks Energy fMRI
Context-FID score (Lower is better) TimeBridge-GP 0.006±.001 0.016±.002 0.067±.007 0.054±.009 0.064±.007 0.096±.008
TimeBridge-GP-$s=0$ 1.956±.149 4.878±.571 6.582±.615 3.531±.263 3.439±.268 10.491±1.066
Correlational score (Lower is better) TimeBridge-GP 0.018±.003 0.173±.025 0.034±.004 0.008±.008 0.902±.218 0.839±.018
TimeBridge-GP-$s=0$ 0.431±.005 0.618±.044 0.441±.010 0.157±.007 3.741±.138 5.235±.091
Discriminative score (Lower is better) TimeBridge-GP 0.002±.002 0.006±.003 0.052±.002 0.049±.014 0.165±.009 0.082±.033
TimeBridge-GP-$s=0$ 0.323±.041 0.445±.018 0.423±.048 0.374±.026 0.499±.000 0.497±.008
Predictive score (Lower is better) TimeBridge-GP 0.093±.000 0.009±.001 0.121±.006 0.037±.000 0.251±.000 0.100±.000
TimeBridge-GP-$s=0$ 0.095±.001 0.018±.002 0.148±.005 0.050±.006 0.255±.002 0.102±.000
Fourier-based loss
Table 22 illustrates the strength of Fourier-based loss in diffusion sampling.
Table 22. Quality comparison of unconditional time series without Fourier.
Metric Methods Synthetic datasets Real datasets
Sines MuJoCo ETTh Stocks Energy fMRI
Context-FID score (Lower is better) TimeBridge-GP 0.006±.001 0.016±.002 0.067±.007 0.054±.009 0.064±.007 0.096±.008
w/o Fourier 0.007±.002 0.018±.003 0.087±.010 0.085±.020 0.093±.011 0.097±.008
Correlational score (Lower is better) TimeBridge-GP 0.018±.003 0.173±.025 0.034±.004 0.008±.008 0.902±.218 0.839±.018
w/o Fourier 0.018±.004 0.172±.022 0.036±.005 0.010±.008 0.942±.198 0.848±.023
Discriminative score (Lower is better) TimeBridge-GP 0.002±.002 0.006±.003 0.052±.002 0.049±.014 0.165±.009 0.082±.033
w/o Fourier 0.004±.004 0.010±.006 0.060±.004 0.035±.027 0.216±.007 0.096±.006
C.7. Visualization
(a) Sines
(b) Mujoco
(c) ETTh
(d) Stocks
(e) Energy
(f) fMRI
Figure 7. Visualization on PCA. Red and blue indicate original and synthetic data, respectively.
Visualization of distribution properties
In addition to visualization of the synthetic data on t-SNE Figure 3, we visualize the synthetic data of TimeBridge using PCA kernel and kernel density estimation in Figures 7 and 8, respectively. More visualization results on baseline methods can be found in the Appendix of Diffusion-TS (Yuan and Qiao, 2024).
(a) Sines
(b) Mujoco
(c) ETTh
(d) Stocks
(e) Energy
(f) fMRI
Figure 8. Visualization of the kernel density estimation. Red and blue indicate original and synthetic data, respectively.
Sampling paths
In Figure 9, we additionally provide the sampling path from prior to data samples (left-to-right), as shown in Figure 4(a). Here, we visualize the priors with GP, linear trend, polynomial trend, and Butterworth trend.
Imputation
In Figure 10, we additionally provide the sampling path from linear spline of masked values as prior to data samples (left-to-right), as shown in Figure 4(b).
In Figure 11, we further visualize the results of imputation, as shown in Figure 5.
(a) Unconditional (GP) - First
(b) Unconditional (GP) - Second
(c) Linear Trend - First
(d) Linear Trend - Second
(e) Polynomial Trend - First
(f) Polynomial Trend - Second
(g) Butterworth Trend - First
(h) Butterworth Trend - Second
Figure 9. Sampling path from (Left) the prior to (Right) the data samples with diverse prior settings on the ETTh dataset.
Figure 10. Sampling path from the prior (Left) to the imputation (Right) of TimeBridge (red) using point-preserving sampling based on observed values (dots) of the truth (green).
(a) Energy
(b) ETTh
Figure 11. Visualization of imputation on (Top) Energy and (Bottom) ETTh datasets. The green dotted lines denote the true values, and the green circles denote the observed values. Diffusion-TS and TimeBridge are visualized with grey and red, along with their interval.